• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name R. Michaels

    Log entry time 18:31:09 on June29,2002

    Entry number 87060

    keyword=more rad measurements, new run plan

    Becky Nevaraz made more radiation measurements at target.
    The levels reduced by only 5 mr in 5 hours and are still too high.
    267 mrad/hr now (what was reported before was slightly wrong).
    Folts cannot do the vacuum coupling work tonight. We will re-measure
    tomorrow at 08:00, but at present the arithmetic says it will be too hot
    all weekend. There were some ideas of using lead blankets, but they
    are not very effective.

    The present plan is the following:

    1)... HAPPEX detector installation now until 23:00.

    2)... At 23:00 (or earlier if installation finished), we do G0 cavity test
    (Ed Strong and R. Suleiman) for 1 hour.

    3)... Owl shift: more parity tests (no target allowed !!) until 06:00.

    4)... 06:00 to 08:00 tune for 5 pass

    5)... 08:00 re-measure radiation near target (Becky Navarez).
    At this point we (inclu. Javier) will decide what to do next.
    If radiation levels still too high we may do HAPPEX detector tests
    (after its installation is complete).